# Periodic orbits and their gravitational waves in EMRIs: supermassive black hole affected by galactic dark matter halos

### v2

@@ Abstract
Periodic orbits exhibiting zoom-whirl behavior have become attractive topics for studying particle dynamics and gravitational wave emission in extreme-mass-ratio inspirals (EMRIs). This study systematically investigates periodic orbits around black holes and their gravitational wave radiation in three dark matter halo environments: NFW, Beta, and Moore models. The dark matter distribution in these models can be effectively incorporated using two parameters—the dark matter characteristic mass and halo characteristic radius. Our results reveal that for a larger dark matter mass and a smaller characteristic radius, the shapes of the periodic orbits and the corresponding gravitational waveforms show more significant deviations from the Schwarzschild case. As the halo characteristic radius increases, the orbital shapes and waveform characteristics gradually converge with the Schwarzschild black hole results. Furthermore, our results also suggest that the NFW and Beta models produce nearly indistinguishable results, while the Moore model shows distinct and resolvable signatures compared with Beta/NFW models. These findings deepen our understanding of dark matter halo effects on periodic motions and gravitational wave signatures, providing guidance on future space-based observation for EMRIs.
I Introduction
@@
In a large number of galaxies, dark matter is known to form a “halo” structure around the supermassive black hole (SMBH) [11, 12]. The mass of these dark matter halos spans multiple orders of magnitude, and their density distribution characteristics are crucial for understanding the formation and evolution of galaxies [13]. Through extensive researches using numerical simulations and astrophysical observations, scholars have developed various theoretical models to describe the density profiles of dark matter halos, such as the Navarro-Frenk-White (NFW) model [14, 15], Beta model [16, 17], Moore model [18], Burkert model [19], Einasto model [20, 21, 22], Dehnen-type Model [23] and Brownstein model [24]. These models have been demonstrated to possess significant value and have practical applications in phenomenological studies, providing critical constraints for both indirect and direct detection of dark matter [25, 26, 27, 28].
@@
The motion of test particles (or small celestial bodies) in a gravitational field serves as a pivotal approach to probe the properties of gravitational fields. Bound precession orbits and circular orbits play a foundational role in the classical tests of gravitational theories, the long-term stability and evolution of gravitational systems (especially strong-field regimes), and gravitational wave emission [29, 30, 31, 32, 33, 34, 35, 36, 37]. However, the motion of small celestial bodies and test particles in strong gravitational fields may also exhibit another important motion pattern—periodic orbits [38, 39, 40, 41]. Periodic orbits have gradually become an inspiring research topic in recent years. They have been extensively studied in quantum-corrected spacetime under quantum gravity theory [42, 43, 44, 45, 46, 47], hairy black holes [48, 49, 50, 51], and other spacetime backgrounds [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. The periodic orbits universally exist in various gravitational systems, and their configuration is typically determined by three integer values: the zoom number $z$, the whirl number $w$, and the vertex number $v$ [38, 39, 41]. Specifically, $z$ denotes the number of “leaves” in a full periodic cycle, with greater $z$ values indicating increased trajectory complexity; $w$ represents the count of additional whirls the particle performs during its outward drift toward the apoapsis; and $v$ characterizes the behavior of the subsequent vertex the particle encounters after leaving the initial vertex (apoapsis). Moreover, the zoom and whirl features of periodic orbits are directly reflected in the gravitational waveform. These periodic orbits and their gravitational-wave radiation signals are promising probes for the exploration of strong gravitational fields. These signals are particularly important for the study of extreme-mass-ratio inspirals (EMRIs) [67, 43, 66]. EMRIs are of significant importance in gravitational wave astronomy. They are composed of a compact small celestial body/test particle that gradually spirals towards the SMBH [68]. They have been identified as promising targets for the next generation of gravitational-wave detectors, such as LISA [69, 70], TianQin [71, 72], Taiji [73, 74] and DECIGO [75].
@@
Recently, theoretical studies have examined the influences of dark matter on the physical processes in the vicinity of SMBHs. These studies involve a wide range of topics and physical processes, including accretion [76, 77], gravitational lensing [78, 79, 80, 81, 82, 83, 84], circular geodesics and black hole shadow  [85, 86, 87, 88, 89, 90, 91, 92], quasi-normal mode and gravitational waves [93, 94, 95, 96, 97, 98]. However, the studies on periodic orbits around SMBH in dark matter environments remain in a very preliminary stage, which are concentrated on a single dark matter model [65, 66]. The comparison of periodic orbits from different dark matter distributions is still absent. The specific impact of dark matter mass and density profiles in halo structures (described by different halo models) on the periodic orbits around SMBHs, as well as the subsequent influence on the generated gravitational wave signals, remain unresolved and challenging issues. These questions are of significant importance for the interpretation of gravitational wave signals in EMRIs in gravitational wave astronomy.
@@
The structure of this paper is as follows: In Section II, we review the spacetime metrics generated by supermassive black holes in dark matter halo environments. We further derive the equations of motion and effective potential for massive test particles, and develop the framework to compute the orbital precession angle $q$ and particle orbits. In Section III, we compare and analyze the periodic orbits of black holes in the backgrounds of different dark matter halos. The influence of dark matter halo parameters on the periodic orbits is emphasised in this section. In Section IV, we explore the gravitational waveforms generated by these periodic orbits. Finally, Section V contains the conclusions and discussions of our work. Throughout the paper, we use the geometric unit system with $G=c=1$.
II Supermassive Black Holes in Dark Matter Halos
@@ II.1 Spacetime Metric
$\mathrm{d}s^{2}=-f(r)\mathrm{d}\mathrm{t}^{2}+f(r)^{-1}\mathrm{d}\mathrm{r}^{2}+r^{2}\mathrm{d}\theta^{2}+r^{2}\mathrm{s}in^{2}\theta\mathrm{d}\phi^{2}.$ (1)
@@
$v_{tg}{}^{2}(r)=\frac{r}{\sqrt{f(r)}}\cdot\frac{d\sqrt{f(r)}}{dr}=\frac{r(\mathrm{dln}\sqrt{f(r)})}{dr}.$ (2)
Based on the rotational velocity relationship in Eq. (2), the dark matter halo contribution to metric function can be obtained by solving the differential equation,
$f_{DM}(r)=\exp\left[2\int\frac{v_{tg}{}^{2}(r)}{r}\mathrm{d}r\right].$ (3)
@@
$M_{DM}(r)=4\pi\int_{0}^{r}\rho(r^{\prime})r^{\prime 2}dr^{\prime},$ (4)
@@
$\rho_{\text{NFW}}(x)$ $=\frac{\rho_{0}}{x(1+x^{2})},$ (5a)
$\rho_{\text{Beta}}(x)$ $=\frac{\rho_{0}}{(1+x^{2})^{3/2}},$ (5b)
$\rho_{\text{Moore}}(x)$ $=\frac{\rho_{0}}{x^{3/2}(1+x^{3/2})},$ (5c)
where $x=r/h$, with $\rho_{0}$ and $h$ representing the characteristic density and characteristic radius (or characteristic scale) of dark matter halo respectively. The effective spacetime metric for supermassive black holes enclosed by dark matter halos is obtained by combining the dark matter density profile $\rho$ and black hole mass $M$ in the Einstein field equations [102]. After some simplification, the metric can be decomposed as $f(r)=f_{\text{DM}}(r)-\frac{2M}{r}$, where $f_{\text{DM}}(r)$ represents the dark matter contribution (see Eq. (3)) and $\frac{2M}{r}$ accounts for the central black hole’s gravitational effect. For different dark matter models, the resulting metric functions take the following forms ((the derivations of metric functions can be found in reference [78]),
$f_{\text{NFW}}(r)=(1+x)^{-\frac{8\pi k}{r}}-\frac{2M}{r},$ (6a)
$f_{\text{Beta}}(r)=e^{-\frac{8\pi k}{r}\sinh^{-1}x}-\frac{2M}{r},$ (6b)
$f_{\mathrm{Moore}}(r)=e^{\frac{16\pi k}{\sqrt{3}h}\arctan\frac{2\sqrt{x}-1}{\sqrt{3}}}\cdot(1+x^{3/2})^{-\frac{16\pi k}{3r}}\cdot\left(\frac{1+x-\sqrt{x}}{1+x+2\sqrt{x}}\right)^{-\frac{8\pi k}{3h}}-\frac{2M}{r}.$ (6c)
@@
II.2 Effective Potential, ISCO and MBO
@@
$p_{\mu}=\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}=g_{\mu\nu}\dot{x}^{\nu},$ (8)
Substituting Eq. (8) into Eq. (1) yields the equations of motion of the particle. For a static and spherically symmetric black hole, the conserved quantities of the system—energy $E$ and angular momentum LL—are related to $p_{t}$ and $p_{\phi}$.
$p_{t}$ $=-f(r)\dot{t}=-E,$ (9a)
$p_{\phi}$ $=r^{2}\sin^{2}\theta\dot{\phi}=L,$ (9b)
$p_{r}$ $=f(r)^{-1}\dot{r},$ (9c)
$p_{\theta}$ $=r^{2}\dot{\theta}.$ (9d)
It can also be derived from Eqs. (9a) and (9b) as follows,
$\dot{t}$ $=\frac{E}{f(r)},$ (10a)
$\dot{\phi}$ $=\frac{L}{r^{2}\sin^{2}\theta}.$ (10b)
Because of the spherical symmetry, we can always constrain the particle orbit to the equatorial plane ($\theta=\pi/2$ and $\dot{\theta}=0$). For a particle following a timelike geodesic in a gravitational field, its four-velocity $\dot{x}^{\mu}$ satisfies the normalization condition,
$g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=-1.$ (11)
Utilizing the relation between conserved quantities and four-velocity in Eqs. (10a), (10b), the normalization of four-velocity Eq. (11) reduces to the following radial orbit equation,
$\dot{r}^{2}+V_{\text{eff}}=E^{2},$ (12)
@@
$V_{\mathrm{eff}}=f(r)\left(1+\frac{L^{2}}{r^{2}}\right)$ (13)
@@
$V_{\text{eff}}=E=1,\quad\frac{dV_{\text{eff}}}{dr}=0.$ (14)
@@
$\dot{r}=0,\quad\frac{dV_{\text{eff}}}{dr}=0,\quad\frac{d^{2}V_{\text{eff}}}{dr^{2}}=0.$ (15)
To investigate the influence of dark matter halos on test particles’ orbital angular momentum, we analyzed the effects of the dark matter mass $k$ on the angular momentum of two critical orbits—MBO and ISCO—using different dark matter halo models, as shown in Fig. 2(a) and 2(b). For both MBO and ISCO orbits, a similar trend is observed: the presence of dark matter halos increases the orbital angular momentum of these orbits. As the dark matter mass parameter $k$ increases, both angular momenta $L_{\text{MBO}}$ and $L_{\text{ISCO}}$ rise accordingly, with the curves for the NFW and Beta models nearly overlapping. For the same dark matter mass parameter $k$, the $L_{\text{MBO}}$ and $L_{\text{ISCO}}$ values of the NFW and Beta models are higher than those of the Moore model. This indicates that under identical dark matter mass and halo scale, the gravitational effects exerted by the NFW and Beta are stronger than those of the Moore.
@@
$L_{\text{ISCO}}\leq L\quad\text{and}\quad E_{\text{ISCO}}\leq E\leq E_{\text{MBO}}=1.$ (16)
@@
II.3 Periodic Orbits and Their Precession Parameter $q$
@@
$q=\frac{\Delta\phi}{2\pi}-1=w+\frac{v}{z}.$ (17)
Here, $\Delta\phi$ denotes the precession angle—the total azimuthal angular change during one complete radial oscillation from periastron $r_{1}$ to apoastron $r_{2}$ and back. The physical significance of the frequency ratio $\frac{\Delta\phi}{2\pi}$ becomes apparent when we consider truly periodic orbits; for such orbits, this ratio must be a rational number [38, 39]. In this expression, the configurations (zwv)(z~w~v) represent the zoom, whirl, and vertex numbers, respectively, which characterize the geometric properties of the periodic orbit. $z$ denotes the number of “leaves” in a full periodic orbit trajectory, with greater $z$ values indicating increased trajectory complexity; $w$ represents the count of additional whirls the particle performs during its outward drift toward the apoapsis; and $v$ characterizes the behavior of the subsequent vertex the particle encounters after leaving the initial vertex (apoapsis) [38, 39, 41]. Two illustrative examples of the orbit with configurations (zwv)=(211)(z~w~v)=(2~1~1) and  (zwv)=(312)(z~w~v)=(3~1~2) are presented in Fig. 3.
Figure 3: Schematic of periodic orbits. (a) The (zwv)=(211)(z~w~v)=(2~1~1) orbit, where z = 2 is visualized by the blue and black curves (representing 2 leaves in one full periodic orbit), and w = 1 is characterized by the purple dotted curve (indicating 1 additional whirl in the inner part of periodic orbit during particles drift outward to apoapsis). (b) The (zwv)(z~w~v) = (3 1 2) orbit illustrating the vertex labeling scheme: vertices are numbered 0, 1, 2 along successive apastra, where v = 2 indicates the orbit skips one vertex and moves to the vertex 2 (at the next apoapsis) after leaving the initial apoapsis at vertex 0.
@@
$\Delta\phi=\oint d\phi=\int_{r_{1}}^{r_{2}}\frac{d\phi}{dr}dr+\int_{r_{2}}^{r_{1}}\frac{d\phi}{dr}dr.$ (18)
To evaluate this integral, we need to express $\frac{d\phi}{dr}$ in terms of the orbital parameters $(E,L)$. By deriving from the equations of motion (Eqs. (10b) and (12)), we obtain,
$\frac{d\phi}{dr}=\frac{d\phi}{d\tau}\cdot\frac{d\tau}{dr}=\pm\frac{L}{r^{2}\sqrt{E^{2}-f(r)\left(1+\frac{L^{2}}{r^{2}}\right)}}.$ (19)
The sign of $\frac{d\phi}{dr}$ depends on the direction of radial motion: it is positive during the outward journey from periastron $r_{1}$ to apoastron $r_{2}$ (where $\dot{r}>0$), and negative during the inward return from $r_{2}$ to $r_{1}$ (where $\dot{r}<0$). Substituting Eq. (19) into Eq. (18) and utilizing the symmetry of the radial motion, the total precession angle becomes,
$\Delta\phi=2\int_{r_{1}}^{r_{2}}\frac{L}{r^{2}\sqrt{E^{2}-f(r)\left(1+\frac{L^{2}}{r^{2}}\right)}}dr.$ (20)
Finally, combining this result with the definition in Eq. (17), we arrive at the feasible integral expression for the precession parameter:
$q=\frac{1}{\pi}\int_{r_{1}}^{r_{2}}\frac{L}{r^{2}\sqrt{E^{2}-f(r)\left(1+\frac{L^{2}}{r^{2}}\right)}}dr-1.$ (21)
The integral formula provides a direct way to compute the precession of periodic orbits. It can be seen that the precession paramater $q$ depends on the energy $E$, angular momentum $L$, and the metric function $f(r)$, which varies with the dark matter halo background surrounding the SMBH, as detailed in Eqs. (6a)–(6c).
@@
$\left(\frac{dr}{d\phi}\right)^{2}=\frac{r^{4}\left[E^{2}-f(r)\left(1+\frac{L^{2}}{r^{2}}\right)\right]}{L^{2}}$ (22)
@@
$\frac{d^{2}r}{d\phi^{2}}=-\frac{r}{2L^{2}}\times\left[r\left(L^{2}\frac{f(r)}{dr}+r^{2}\frac{f(r)}{dr}-4E^{2}r\right)+2f(r)\left(L^{2}+2r^{2}\right)\right]$ (23)

### v3

@@ Abstract
Periodic orbits exhibiting zoom-whirl behavior have become attractive topics for studying particle dynamics and gravitational wave emission in extreme-mass-ratio inspirals (EMRIs). This study systematically investigates periodic orbits around black holes and their gravitational wave radiation in three dark matter halo environments: NFW, Beta, and Moore models. The dark matter distribution in these models can be effectively incorporated using two parameters — the dark matter characteristic mass and halo characteristic radius. Our results reveal that for a larger dark matter mass and a smaller characteristic radius, the shapes of the periodic orbits and the corresponding gravitational waveforms show more significant deviations from the Schwarzschild case. As the halo characteristic radius increases, the orbital shapes and waveform characteristics gradually converge with the Schwarzschild black hole results. Furthermore, our results also suggest that the NFW and Beta models produce nearly indistinguishable results, while the Moore model shows distinct signatures compared with Beta/NFW models. These findings deepen our understanding of dark matter halo effects on periodic motions and gravitational wave signatures.
Keywords: GR black holes; Gravitational waves in GR and beyond: theory;
dark matter theory; astrophysical black holes
1 Introduction
@@
In a large number of galaxies, dark matter is known to form a halo structure around the supermassive black hole (SMBH) [11, 12]. The mass of these dark matter halos spans multiple orders of magnitude, and their density distribution characteristics are crucial for understanding the formation and evolution of galaxies [13]. Through extensive researches using numerical simulations and astrophysical observations, scholars have developed various theoretical models to describe the density profiles of dark matter halos, such as the Navarro-Frenk-White (NFW) model [14, 15], Beta model [16, 17], Moore model [18], Burkert model [19], Einasto model [20, 21, 22], Dehnen-type Model [23] and Brownstein model [24]. These models have been demonstrated to possess significant value and have practical applications in phenomenological studies, providing critical constraints for both indirect and direct detection of dark matter [25, 26, 27, 28].
@@
The motion of test particles (or small celestial bodies) in a gravitational field serves as a pivotal approach to probe the properties of gravitational fields. Bound precession orbits and circular orbits play a foundational role in the classical tests of gravitational theories, the long-term stability and evolution of gravitational systems (especially strong-field regimes), and gravitational wave emission [29, 30, 31, 32, 33, 34, 35, 36, 37]. However, the motion of small celestial bodies and test particles in strong gravitational fields may also exhibit another important motion pattern—periodic orbits [38, 39, 40, 41]. Periodic orbits have gradually become an inspiring research topic in recent years. They have been extensively studied in quantum-corrected spacetime under quantum gravity theory [42, 43, 44, 45, 46, 47], hairy black holes [48, 49, 50, 51], and other spacetime backgrounds [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. The periodic orbits universally exist in various gravitational systems, and their configuration is typically determined by three integer values: the zoom number $z$, the whirl number $w$, and the vertex number $v$ [38, 39, 40]. Specifically, $z$ denotes the number of leaves in a full periodic period, with greater $z$ values indicating increased trajectory complexity; $w$ represents the count of additional whirls the particle performs during its outward drift toward the apoapsis; and $v$ characterizes the behavior of the subsequent vertex the particle encounters after leaving the initial vertex (apoapsis). Moreover, the zoom and whirl features of periodic orbits are directly reflected in the gravitational waveform. These periodic orbits and their gravitational-wave radiation signals are promising probes for the exploration of strong gravitational fields. These signals are particularly important for the study of extreme-mass-ratio inspirals (EMRIs) [68, 43, 67]. EMRIs are of significant importance in gravitational wave astronomy. They are composed of a compact small celestial body/test particle that gradually spirals towards the SMBH [69]. They have been identified as promising targets for the next generation of gravitational-wave detectors, such as LISA [70, 71], TianQin [72, 73], Taiji [74, 75] and DECIGO [76].
@@
Recently, theoretical studies have examined the influences of dark matter on the physical processes in the vicinity of SMBHs. These studies involve a wide range of topics and physical processes, including accretion [77, 78], gravitational lensing [79, 80, 81, 82, 83, 84, 85], circular geodesics and black hole shadow  [86, 87, 88, 89, 90, 91, 92, 93], quasi-normal mode and gravitational waves [94, 95, 96, 97, 98, 99, 100]. However, the studies on periodic orbits around SMBH in dark matter environments remain in a very preliminary stage, which are concentrated on a single dark matter model [65, 67, 66]. The comparison of periodic orbits from different dark matter distributions is still absent. The specific impact of dark matter mass and density profiles in halo structures (described by different halo models) on the periodic orbits around SMBHs, as well as the subsequent influence on the generated gravitational wave signals, remain unresolved and challenging issues. These questions are of significant importance for the interpretation of gravitational wave signals in EMRIs in gravitational wave astronomy.
@@
The structure of this paper is as follows: In Section 2, we review the spacetime metrics generated by supermassive black holes in dark matter halo environments. We further derive the equations of motion and effective potential for massive test particles, and develop the framework to compute the orbital precession angle $q$ and particle orbits. In Section 3, we compare and analyze the periodic orbits of black holes in the backgrounds of different dark matter halos. The influence of dark matter halo parameters on the periodic orbits is emphasised in this section. In Section 4, we explore the gravitational waveforms generated by these periodic orbits. Finally, Section 5 contains the conclusions and discussions of our work. Throughout the paper, we use the geometric unit system with $G=c=1$.
2 Supermassive Black Holes in Dark Matter Halos
@@ 2.1 Spacetime Metric
$\mathrm{d}s^{2}=-f(r)\mathrm{d}\mathrm{t}^{2}+f(r)^{-1}\mathrm{d}\mathrm{r}^{2}+r^{2}\mathrm{d}\theta^{2}+r^{2}\mathrm{s}in^{2}\theta\mathrm{d}\phi^{2}.$ (2.1)
@@
$v_{tg}{}^{2}(r)=\frac{r}{\sqrt{f(r)}}\cdot\frac{d\sqrt{f(r)}}{dr}=\frac{r(\mathrm{dln}\sqrt{f(r)})}{dr}.$ (2.2)
Based on the rotational velocity relationship in Eq. (2.2), the dark matter halo contribution to metric function can be obtained by solving the differential equation,
$f_{DM}(r)=\exp\left[2\int\frac{v_{tg}{}^{2}(r)}{r}\mathrm{d}r\right].$ (2.3)
@@
$M_{DM}(r)=4\pi\int_{0}^{r}\rho(r^{\prime})r^{\prime 2}dr^{\prime},$ (2.4)
@@
$\rho_{\text{NFW}}(x)$ $=\frac{\rho_{0}}{x(1+x^{2})},$ (2.5a)
$\rho_{\text{Beta}}(x)$ $=\frac{\rho_{0}}{(1+x^{2})^{3/2}},$ (2.5b)
$\rho_{\text{Moore}}(x)$ $=\frac{\rho_{0}}{x^{3/2}(1+x^{3/2})},$ (2.5c)
where $x=r/h$, with $\rho_{0}$ and $h$ representing the characteristic density and characteristic radius (or characteristic scale) of dark matter halo respectively. The effective spacetime metric for supermassive black holes enclosed by dark matter halos is obtained by combining the dark matter density profile $\rho$ and black hole mass $M$ in the Einstein field equations [104]. After some simplification, the metric can be decomposed as $f(r)=f_{\text{DM}}(r)-\frac{2M}{r}$, where $f_{\text{DM}}(r)$ represents the dark matter contribution (see Eq. (2.3)) and $\frac{2M}{r}$ accounts for the central black hole’s gravitational effect. For different dark matter models, the resulting metric functions take the following forms ((the derivations of metric functions can be found in reference [79]),
$f_{\text{NFW}}(r)=(1+x)^{-\frac{8\pi k}{r}}-\frac{2M}{r},$ (2.6a)
$f_{\text{Beta}}(r)=e^{-\frac{8\pi k}{r}\sinh^{-1}x}-\frac{2M}{r},$ (2.6b)
$f_{\mathrm{Moore}}(r)=e^{\frac{16\pi k}{\sqrt{3}h}\arctan\frac{2\sqrt{x}-1}{\sqrt{3}}}\cdot(1+x^{3/2})^{-\frac{16\pi k}{3r}}\cdot\left(\frac{1+x-\sqrt{x}}{1+x+2\sqrt{x}}\right)^{-\frac{8\pi k}{3h}}-\frac{2M}{r}.$ (2.6c)
@@
2.2 Effective Potential, ISCO and MBO
@@
$p_{\mu}=\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}}=g_{\mu\nu}\dot{x}^{\nu},$ (2.8)
Substituting Eq. (2.8) into Eq. (2.1) yields the equations of motion of the particle. For a static and spherically symmetric black hole, the conserved quantities of the system — energy $E$ and angular momentum $L$ — are related to $p_{t}$ and $p_{\phi}$.
$p_{t}$ $=-f(r)\dot{t}=-E,$ (2.9a)
$p_{\phi}$ $=r^{2}\sin^{2}\theta\dot{\phi}=L,$ (2.9b)
$p_{r}$ $=f(r)^{-1}\dot{r},$ (2.9c)
$p_{\theta}$ $=r^{2}\dot{\theta}.$ (2.9d)
It can also be derived from Eqs. (2.9a) and (2.9b) as follows,
$\dot{t}$ $=\frac{E}{f(r)},$ (2.10a)
$\dot{\phi}$ $=\frac{L}{r^{2}\sin^{2}\theta}.$ (2.10b)
[…]
$g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=-1.$ (2.11)
Utilizing the relation between conserved quantities and four-velocity in Eqs. (2.10a), (2.10b), the normalization of four-velocity Eq. (2.11) reduces to the following radial orbit equation,
$\dot{r}^{2}+V_{\text{eff}}=E^{2},$ (2.12)
@@
$V_{\mathrm{eff}}=f(r)\left(1+\frac{L^{2}}{r^{2}}\right)$ (2.13)
@@
$V_{\text{eff}}=E=1,\quad\frac{dV_{\text{eff}}}{dr}=0.$ (2.14)
@@
$\dot{r}=0,\quad\frac{dV_{\text{eff}}}{dr}=0,\quad\frac{d^{2}V_{\text{eff}}}{dr^{2}}=0.$ (2.15)
To investigate the influence of dark matter halos on test particles’ orbital angular momentum, we analyzed the effects of the dark matter mass $k$ on the angular momentum of two critical orbits — MBO and ISCO — using different dark matter halo models,ăasăshown in Fig. 2(a) and 2(b). For both MBO and ISCO orbits, a similar trend is observed: the presence of dark matter halos increases the orbital angular momentum of these orbits. As the dark matter mass parameter $k$ increases, both angular momenta $L_{\text{MBO}}$ and $L_{\text{ISCO}}$ rise accordingly, with the curves for the NFW and Beta models nearly overlapping. For the same dark matter mass parameter $k$, the $L_{\text{MBO}}$ and $L_{\text{ISCO}}$ values of the NFW and Beta models are higher than those of the Moore model. This indicates that under identical dark matter mass and halo scale, the gravitational effects exerted by the NFW and Beta are stronger than those of the Moore.
@@
$L_{\text{ISCO}}\leq L\quad\text{and}\quad E_{\text{ISCO}}\leq E\leq E_{\text{MBO}}=1.$ (2.16)
@@
2.3 Periodic Orbits and Their Precession Parameter $q$
@@
$q=\frac{\Delta\phi}{2\pi}-1=w+\frac{v}{z}.$ (2.17)
Here, $\Delta\phi$ denotes the precession angle — the total azimuthal angular change during one complete radial oscillation from periastron $r_{1}$ to apoastron $r_{2}$ and back. The physical significance of the frequency ratio $\frac{\Delta\phi}{2\pi}$ becomes apparent when we consider truly periodic orbits; for such orbits, this ratio must be a rational number [38, 39]. In this expression, the configurations (zwv)(z~w~v) represent the zoom, whirl, and vertex numbers, respectively, which characterize the geometric properties of the periodic orbit. $z$ denotes the number of leaves in a full periodic orbit trajectory, with greater $z$ values indicating increased trajectory complexity; $w$ represents the count of additional whirls the particle performs during its outward drift toward the apoapsis; and $v$ characterizes the behavior of the subsequent vertex the particle encounters after leaving the initial vertex (apoapsis) [38, 39, 40]. Two illustrative examples of the orbit with configurations (zwv)=(211)(z~w~v)=(2~1~1) and (zwv)=(312)(z~w~v)=(3~1~2) are presented in Fig. 3.
Figure 3: Schematic of periodic orbits. (a) The (zwv)=(211)(z~w~v)=(2~1~1) orbit, where z = 2 is visualized by the blue and black curves (representing 2 leaves in one full orbital period), and w = 1 is characterized by the purple dotted curve (indicating 1 additional whirl in the inner part of periodic orbit during particles drift outward to apoapsis). (b) The (zwv)(z~w~v) = (3 1 2) orbit illustrating the vertex labeling scheme: vertices are numbered 0, 1, 2 along successive apastra, where v = 2 indicates the orbit skips one vertex and moves to the vertex 2 (at the next apoapsis) after leaving the initial apoapsis at vertex 0.
@@
$\Delta\phi=\oint d\phi=\int_{r_{1}}^{r_{2}}\frac{d\phi}{dr}dr+\int_{r_{2}}^{r_{1}}\frac{d\phi}{dr}dr.$ (2.18)
To evaluate this integral, we need to express $\frac{d\phi}{dr}$ in terms of the orbital parameters $(E,L)$. By deriving from the equations of motion (Eqs. (2.10b) and (2.12)), we obtain,
$\frac{d\phi}{dr}=\frac{d\phi}{d\tau}\cdot\frac{d\tau}{dr}=\pm\frac{L}{r^{2}\sqrt{E^{2}-f(r)\left(1+\frac{L^{2}}{r^{2}}\right)}}.$ (2.19)
The sign of $\frac{d\phi}{dr}$ depends on the direction of radial motion: it is positive during the outward journey from periastron $r_{1}$ to apoastron $r_{2}$ (where $\dot{r}>0$), and negative during the inward return from $r_{2}$ to $r_{1}$ (where $\dot{r}<0$). Substituting Eq. (2.19) into Eq. (2.18) and utilizing the symmetry of the radial motion, the total precession angle becomes,
$\Delta\phi=2\int_{r_{1}}^{r_{2}}\frac{L}{r^{2}\sqrt{E^{2}-f(r)\left(1+\frac{L^{2}}{r^{2}}\right)}}dr.$ (2.20)
Finally, combining this result with the definition in Eq. (2.17), we arrive at the feasible integral expression for the precession parameter:
$q=\frac{1}{\pi}\int_{r_{1}}^{r_{2}}\frac{L}{r^{2}\sqrt{E^{2}-f(r)\left(1+\frac{L^{2}}{r^{2}}\right)}}dr-1.$ (2.21)
The integral formula provides a direct way to compute the precession of periodic orbits. It can be seen that the precession paramater $q$ depends on the energy $E$, angular momentum $L$, and the metric function $f(r)$, which varies with the dark matter halo background surrounding the SMBH, as detailed in Eqs. (2.6a)-(2.6c).
@@
$\left(\frac{dr}{d\phi}\right)^{2}=\frac{r^{4}\left[E^{2}-f(r)\left(1+\frac{L^{2}}{r^{2}}\right)\right]}{L^{2}}$ (2.22)
@@
$\frac{d^{2}r}{d\phi^{2}}=-\frac{r}{2L^{2}}\times\left[r\left(L^{2}\frac{f(r)}{dr}+r^{2}\frac{f(r)}{dr}-4E^{2}r\right)+2f(r)\left(L^{2}+2r^{2}\right)\right]$ (2.23)